an. 1 Retired a piece of machinery that was purchased on January 1, 2009. The machine cost $62,900 on that date. It had a useful life of 10 years with no salvage value. computer that was purchased on January 1, 2016. The computer cost $43,200. It had a useful life of 5 years with no salvage value. The computer was sold for in the order presented in the problem. If no entry is required, select No Entry for the account titles and enter 0 for the amounts. Do not round
media%2Ffd0%2Ffd016c45-967d-4712-94e9-c0
Answer #1
Date Account Description Debit Credit
1-Jan Accumulated Depreciation- Machinery $        62,900
Machinery $        62,900
(To record the ritirement of machinery)
30-Jun Depreciation $          4,320
Accumulated Depreciation- Computer $          4,320
(To Record the Depreciation till Disposal date)
30-Jun Cash $        15,100
Accumulated Depreciation (4320+25920) $        30,240
Computer $        43,200
Gain on Disposal $          2,140
(To Record Disposal of Computer)
31-Dec Depreciation $          5,600
Accumulated Depreciation $          5,600
(To Record Depreciation on Truck)
31-Dec Loss on Retirement $          8,600
Accumulated Depreciation $        28,000
Truck $        36,600
(To Record the Discarding of Truck)
Cost of Computer = $ 43,200
Useful Life = 5 Years
Depreciation per Year = $    8,640
Accumulated Depreciation on 1 Jan 2019(8640*3) = $ 25,920
Depreciation- 2019 ( 8640/12*6) = $    4,320
Cost of Truck = $ 36,600
Salvage Value = $    3,000
Depreciable Value = $ 33,600
Useful Life = 6 Years
Depreciation per Year = $    5,600
Accumulated Depreciation on 1 Jan 2019(5600*4) = $ 22,400
Depreciation- 2019 = $    5,600
Total Accumulated Depreciation = $ 28,000

Please note in the last Journal entry I have assumed that the Truck is discarded with no cash flow in the absence of information.


Dear Student,

Best effort has been made to give quality and correct answer. But if you find any issues please comment your concern. I will definitely resolve your query.